ABC Company employs a periodic inventory system and sells its inventory to customers for $27 per unit. ABC Company had the following inventory information available for the month of May:

May 1 Beginning inventory 2,400 units @ $14 cost per unit

May 8 Sold 1,000 units

May 13 Purchased 1,300 units @ $17 cost per unit

May 18 Sold 1,200 units

May 21 Purchased 1,100 units @ $11 cost per unit

May 28 Sold 300 units

May 30 Purchased 700 units @ $10 cost per unit

During May, ABC Company reported operating expenses of $14,500 and had an income tax rate of 43%. Calculate the amount of net income shown on ABC Company's income statement for May using the LIFO method.
Answer #1

LIFO method states that goods purchased later should be sold first

Records are maintained and updates at the end of the period under periodic method

Total units sold = 1000+1200+300= 2500 units

Cost of goods sold = 700*10 + 1100*11+700*17

= $31,000

Net Income = (Sales Revenue – Cost of Goods sold – Operating expenses)(1-Tax rate)

= (2500*27 – 31000 – 14500)*(1-43%)

= $12,540